OpenStudy (anonymous):

How many grams of iron metal do you expect to be produced when 325 grams of an 87.5 percent by mass iron (II) nitrate solution reacts with excess aluminum metal? Show all of the work needed to solve this problem. 2 Al (s) + 3 Fe(NO3)2 (aq) 3 Fe (s) + 2 Al(NO3)2 (aq)

OpenStudy (chmvijay):

what is molar mass of Fe(NO3)2

OpenStudy (anonymous):

180

OpenStudy (chmvijay):

and atomic mass of Fe

OpenStudy (anonymous):

55.845

OpenStudy (chmvijay):

180*3 gram of iron nitrate is used to produce the 55.845*3 g of Fe Right

OpenStudy (anonymous):

yeah

OpenStudy (chmvijay):

540 gram of iron nitrate is used to produce the 167.53 g of Fe 325 gram of iron nitrate is used to produces = 167.53*325 /540 = can u answer this

OpenStudy (anonymous):

100.82 ?

OpenStudy (chmvijay):

this value is when 100 % of Fe nitrate we considered as to be reacted but they say around 87.5 wt % only reacted then 100 % - 100.82 87.5 % = 100.82 *87.5 /100 = find it did you get it

OpenStudy (anonymous):

88.21

OpenStudy (chmvijay):

yaaa hats the answer

OpenStudy (anonymous):

thank you! @chmvijay

OpenStudy (chmvijay):

you are welcome dude Just enjoy the chemistry with LOVE:)

OpenStudy (anonymous):

should i roung up to 88.3?

OpenStudy (chmvijay):

no u got 88.21 right u cant do to 88.3 u can do only 88.2 why do u want it to be 88.3

OpenStudy (anonymous):

ok just wondering thanks!